Prove that

(

)

(

)

1,

1

,

nPnnPnn

 − − =

.



1
Expert's answer
2021-10-18T15:22:35-0400

Prove that "^nP_{n}=^nP_{n-1}"


Proof:

"^nP_{n}=n!"


"^nP_{n-1}=\frac{n!}{[n-(n-1)]!}=\frac{n!}{1}=n!"